2x = 1/4 the x is squared
7x=10 the x is squared
8x=6.5 the x is squared
3x=4 the x is squared


Answer by Alan3354(69443)

2x = 1/4 the x is squared
Do you mean 2x^2 = 1/4?
Use ^ (Shift 6) for exponents.
If you can't do that, stop posting these problems.
========================
7x=10 the x is squared
8x=6.5 the x is squared
3x=4 the x is squared